The moon has a mass of 7.32 Γ— 1022 π‘˜π‘” and a radius of 1,609.4 π‘˜π‘š. Calculate the value of "𝑔" at the surface of the moon.


1
Expert's answer
2021-12-15T11:55:33-0500

"g=F/m=\frac{GMm}{r^2m}=\frac{GM}{r^2}=\frac{6.67\cdot10^{-11}\cdot7.32\cdot10^{22}}{1609400^2}=1.88\ (m/s^2)"